A plastic light pipe has an index of refraction of 1.66. For total internal reflection, what is the minimum angle of incidence if the pipe is in the following media?

a) air?=_____ degrees

b) water?=____ degrees
Answer #1

a)
n1=1.66
refractive indix of air=n2=1
critical angle= asin (n2/n1)
             =asin (1/1.66)
            =37 degree
a)
n1=1.66
refractive indix of water=n2=4/3=1.33
critical angle= asin (n2/n1)
             =asin (1.33/1.66)
            =53.4 degree
Answer #2

a)

in air

1.66*sinC =1*sin90

critical angle C =sin-1 (1/1.66) =37 degrres

b)

in water

1.66sinC =1.33sin90

C=sin-1 (1.33/1.66) =53.24 degrees